How do i find the 15th term of the series 3,7,11,15...
Answer by jim_thompson5910(35256)

nth term:

an = a1 + (n-1)*d

an = 3 + (n-1)*4

an = 3 + 4n - 4

an = 4n - 1

-------------------
15th term

an = 4n - 1

a15 = 4(15) - 1

a15 = 60 - 1

a15 = 59